# When Location Shapes Choice:

### v3

@@ 3.2 Hardness of Placement
Recall that the cardinality constrained assortment optimization problem is a special case of Placement (Section 2.1). In this subsection, we argue that, unless P=NP, a $(1-1/e)$ approximation is the best achievable guarantee in polynomial time for cardinality constrained assortment optimization under a Mixture of Multinomial Logit choice model (MMNL) and uniformly priced products. This proves that the $(1-1/e)$-approximation guarantee for the Placement problem with uniformly priced products (derived in the previous section) is best possible.
@@
The proof of this theorem is included in Appendix 8.3.
@@ 4.1 Warm-up: Placement on a Line
Proof 4.2
Proof. Fix $k\in[m]$. First, let us show that the optimal placement of $S^{*}_{k}$ is at most $\frac{O(1)}{k}\cdot\textsc{OPT}$ in the worst case. Consider the following instance on a line and choice given by MNL model. We use $v_{i}$ to denote the preference weight of product $i$ under the MNL model. We have $m$ locations in the line: $\{1,2,\ldots,m\}$ with $1$ being the topmost. Recall $\theta_{j}$ is the probability of visiting the subset of locations $(1:j)$. Let $\theta_{1}=1$ and $\theta_{j}=0$ for $j\in\{2,\ldots,m\}$. Let $N$ be a universe of $k+1$ distinct products with $N=U\cup\{q\}$. There are $k$ products in $U$ where for all $i\in U$, $r_{i}=k$ and $v_{i}=1/k$. Additionally, $r_{q}=k/2$ and $v_{q}=1$. Under MNL, we can compute optimally cardinality-constrained assortments. So, $S^{*}_{k}$ here denotes the optimal assortment of size $k$ under MNL. First, let us show that $S^{*}_{k}=U$. We have $R(U)=k/2$. Moreover, let $p<k$ and let $S\subseteq U$ such that $S$ has $p$ products. Since all products in $U$ are similar, it does not matter which $p$ products we pick. We have $R(S\cup\{q\})=\frac{p+k/2}{1+p/k+1}$, which, for $p<k$, is increasing in $p$. This implies that $R(S\cup\{q\})<k/2$. We can also verify that $R(S)<k/2$ for any $p$. Therefore, we have that $S_{k}^{*}=U$. Now consider $X_{1}$ a placement of $U$ on the line. Note that all locations have 00 visiting probability except the first one. Since, all products in $U$ are similar, it doesn’t matter which product we place in location $1$, we get the following revenue
$\mathcal{R}(X_{1})=\frac{1}{1+1/k}=\frac{k}{k+1}.$
On the other hand, denote $X_{2}$ a placement of product $q$ in the first location. The rest of locations are empty. We have
$\mathcal{R}(X_{2})=R(\{q\})=\frac{k/2}{1+1}=\frac{k}{4}.$
Therefore, for any placement $X_{1}$, we have $\mathcal{R}(X_{1})=\frac{O(1)}{k}\cdot\mathcal{R}(X_{2})\leq\frac{O(1)}{k}\cdot\textsc{OPT}$.
Second, let us show that the optimal placement of $S^{*}_{k}$ is at most $O(1)\cdot\frac{k}{m}\cdot\textsc{OPT}$ in the worst case. Consider the following instance on a line and choice given by MNL model. Let $\theta_{j}=1/m$ be the probability of visiting $(1:j)$. Consider $N$ a universe of $m$ distinct products such that for all $i\in N$, $r_{i}=m$ and $v_{i}=1/m$. Clearly $S_{k}^{*}=\{1,\ldots,k\}$ and $S_{m}^{*}=N$. We have $R(S_{k}^{*})=k/(1+k/m)$ and $R(S_{m}^{*})=m/2$. Let $X_{1}$ be a placement of products $\{1,\ldots,k\}$ on the line. The expected revenue of placement $X_{1}$ is given by
$\mathcal{R}(X_{1})=\sum_{j=1}^{m}\theta_{j}R(X_{1}(1:j))=\sum_{j=1}^{k}\theta_{j}R(X_{1}(1:j))+\sum_{j=k+1}^{m}\theta_{j}R(X_{1}(1:j)).$
We have for any $j\leq k$, $R(X_{1}(1:j))\leq R(\{1,\ldots,j\})=j/(1+j/m)\leq j$. We have for any $j\geq k+1$, $R(X_{1}(1:j))\leq R(\{1,\ldots,k\})=k/(1+k/m)\leq k$. Therefore,
$\mathcal{R}(X_{1})\leq\frac{1}{m}\left(\sum_{j=1}^{k}j+\sum_{j=k+1}^{m}k\right)=\frac{k(k+1)}{2m}+\frac{(m-k)k}{m}=O(1)k.$
On the other hand, by considering $X_{2}$, the placement where product $j\in N$ is placed at location $j$ for every $j\in N$, we get
$\mathcal{R}(X_{2})=\sum_{j=1}^{m}\theta_{j}R(\{1,\ldots,j\})=1/m\sum_{j=1}^{m}\frac{j}{1+j/m}\geq 1/m\sum_{j=1}^{m}\frac{j}{2}=\frac{m+1}{4}.$
Therefore, for any placement $X_{1}$, we have $\mathcal{R}(X_{1})=O(1)\cdot\frac{k}{m}\cdot\mathcal{R}(X_{2})\leq O(1)\cdot\frac{k}{m}\cdot\textsc{OPT}$.
@@
From Lemma 4.1, we observe that selecting products solely from a fixed $S_{k}^{*}$ can be highly restrictive. In the worst case, the expected revenue from such a placement does not exceed $\frac{O(1)}{\sqrt{m}}\cdot\textsc{OPT}$. Motivated by this lemma, we explore algorithms that, for each $k\in[m]$, compute a placement using products from $S_{k}^{*}$ and then select the best among these placements. We demonstrate that this approach markedly enhances the solution to the Placement on a Line problem.
Importance of product selection from the Best of Many $\bf S_{k}^{*}$. For each $k\in[m]$, let $X^{*}_{k}$ represent the placement that positions products from $S^{*}_{k}$ in the first $k$ locations of the line (in any given order) while leaving the subsequent locations empty. Recall from Section 2, that we can always make this placement feasible without loss of generality by filling empty locations with the highest price product. Define $Y$ as the most optimal placement from the set $\{X^{*}_{k}\}_{k\in[m]}$. Specifically, consider the placement solution for the problem given by:
[…]
In the next lemma, we establish that $Y$ gives $\frac{\alpha}{\log\rho}$ approximation to Placement on a Line. The result holds for all choice models satisfying Assumption 2, including multi-purchase models.
Recall that, $j_{s}$ ($\geq 1$) and $j_{l}$ denote the size of the smallest (non-empty) set and largest set (respectively) with non-zero support in the browsing distribution. Also, $\rho=\frac{j_{l}+1}{j_{s}}=2^{d},$ for some integer $d\in\{1,2,\cdots,\lceil\log_{2}(m+1)\rceil\}$.
@@
Let $O:G\to N$ denote an optimal placement for Placement on a Line and let $\mathcal{R}(O)=\textsc{OPT}$. Recall that $O(1:k)$ denotes the set of products placed in locations $\{1,\cdots,k\}$. For $\ell\geq 1$, let $\mathcal{R}_{\ell}(O)$ denote the revenue contribution from products in location $2^{\ell-1}j_{s}$ to $12^{\ell}j_{s}-$ 1, i.e.,
$\mathcal{R}_{\ell}(O)=\sum_{j=2^{\ell-1}j_{s}}^{2^{\ell}j_{s}-1}\left[r_{O(j)}\,\sum_{k\geq j}\theta_{k}\,\phi\left(O(j),O(1:k)\right)\right].$ (2)
@@
[…]
We have $G_{\ell,k}\subseteq(2^{\ell-1}j_{s}:2^{\ell}j_{s}-1)$. Therefore, by using (2) and the weak-rationality of $\phi$ (Assumption 2), we get
$\mathcal{R}_{\ell}(O)\leq\sum_{j=2^{\ell-1}j_{s}}^{2^{\ell}j_{s}-1}r_{O(j)}\sum_{k=j}^{j_{l}}\theta_{k}\phi\left(O(j),O(G_{\ell,k})\right).$
Observe that the RHS is the expected total revenue of the placement solution where we place products $O(2^{\ell-1}j_{s}:2^{\ell}j_{s}-1)$ between $2^{\ell-1}j_{s}$ and $12^{\ell}j_{s}-$ 1, same as $O$, and keep all other positions vacant. Let $O_{\ell}$ denote this solution. Recall that we can relax the definition of a placement to allow vacant locations, without loss of generality, as explained in Remark 2.1 (see Section 2). We have
Consider some $\ell$ in $[d]$. The total revenue of placement $X^{*}_{2^{\ell-1}j_{s}}$ can be lower bounded by the revenue of placement $O_{\ell}$ as follows,
$\mathcal{R}(O_{\ell})$ $=$ $\sum_{j=2^{\ell-1}j_{s}}^{2^{\ell}j_{s}-1}r_{O(j)}\sum_{k=j}^{j_{l}}\theta_{k}\phi\left(O(j),O(G_{\ell,k})\right),$
$=$ $\sum_{k=2^{\ell-1}j_{s}}^{j_{l}}\theta_{k}\,R(O(G_{\ell,k})),$
$\leq$ $\frac{1}{\alpha}\sum_{k=2^{\ell-1}j_{s}}^{j_{l}}\theta_{k}\,R\left(S^{*}_{2^{\ell-1}j_{s}}\right),$
$\leq$ $\frac{1}{\alpha}\mathcal{R}(X^{*}_{2^{\ell-1}j_{s}}),$
where the first inequality follows from (1) and the fact that $|O(G_{\ell,k})|\leq|G_{\ell,k}|\leq 2^{\ell-1}j_{s}$. Thus,
$\mathcal{R}(O)\leq\frac{1}{\alpha}\sum_{\ell=1}^{d}\mathcal{R}(X^{*}_{2^{\ell-1}j_{s}})\leq\frac{\log_{2}(\rho)}{\alpha}\,\mathcal{R}(Y),$
[…]
Remarks: Recall that in the proof of Lemma 4.3, we make use of at most $\log_{2}\rho$ candidate placements. In Section 4.3, we show that our analysis is tight even if we take the best placement out of the $m$ candidates $\{X^{*}_{k}\}_{k\in[m]}$. Different methods for placing/ordering the products from sets $S^{*}_{k}$ do not improve the worst case guarantee.
@@ 4.2 Main Results: Approximation Guarantee
Given an assortment, finding the optimal placement for general browsing distributions is significantly more involved as customers may start browsing from various locations and may also travel between locations in myriad ways. In particular, the optimal solution may now require copies of a product to be placed in multiple locations. We illustrate the importance of repeating products with a simple example.
@@
denote the final (randomized) placement solution.
For any given $k\in[m]$, the random placement $X^{r}_{k}$ is oblivious to how customers travel between locations. In contrast, for the Placement on a Line problem, we greedily placed the products in $S^{*}_{k}$ at the top $k$ spots. The oblivious nature of randomized placement is pivotal in proving analytical guarantees.
To analyze our algorithm, we use a method akin to the probabilistic method to show that the algorithm performs “well” in expectation. Standard repetition can be used to obtain an algorithm that generates a good placement with high probability. Later, we also give a derandomized algorithm with a similar guarantee. We now introduce an equivalent (but perhaps somewhat less intuitive) description of the browsing distribution that will be useful in the analysis. Let $\pi:[m]\to[m]$ denote a permutation over $[m]$ and let $\Pi$ denote the set of all permutations over $[m]$. Let $\theta^{\pi}_{j}$ denote the probability that customer visits locations $\{\pi(1),\pi(2),\cdots,\pi(j)\}$ in $G$ (in order). Let $L=\{\pi(1),\pi(2),\cdots,\pi(j)\}$. Recall that, $\mathbb{P}_{B}(L)$ represents the probability that $L$ is the set of locations visited by the customer in any order. In contrast, $\theta^{\pi}_{j}$ is the probability that customer visits the locations in $L$ in order. As we demonstrate later, this refinement is useful for the technical development of our ideas but is not needed for implementing our algorithms. Let OPT represent the optimal value of the Placement problem. Let $X^{\pi}(1:j)$ denote the set of products placed at locations $\{\pi(1),\cdots,\pi(j)\}$. Our goal is to find placement $X$ that maximizes the total expected revenue,
@@
To prove Theorem 4.6, we first show the result for a line graph (Lemma 4.9). We then use the result for a line graph as a building block for establishing Theorem 4.6. Recall that in Lemma 4.3 we showed that the best (deterministic) placement from the set $\{X^{*}_{k}\}_{k\in[m]}$, is at least $\frac{\alpha}{{\log_{2}\rho}}$-approximate. Also, recall that, the exact order of placement had no impact on the result, i.e., $X^{*}_{k}$ can be any placement where the products in $S^{*}_{k}$ are placed in the first $k$ spots on the line. As an immediate consequence, we have that a random placement of products in $S^{*}_{k}$ in the first $k$ spots is $\frac{\alpha}{{\log_{2}\rho}}$-approximate on every sample path. However, this does not give an approximation guarantee for the randomized placement $Y$ output by our algorithm. The difference being that in our algorithm we now allow repetition of products and the candidate placement $X^{r}_{k}$ may not place some products in $S^{*}_{k}$ at any location. In fact, the value of $\mathcal{R}(X^{r}_{k})$ may be very low with some probability, such as, when a “low revenue” product is placed at every location. The following lemma shows that such possibilities have a limited effect on the expected revenue.
Similar to the proof of Lemma 4.3, recall $\rho=\frac{j_{l}+1}{j_{s}}=2^{d}$ for some integer $d\in\{1,2,\cdots,{\log_{2}(m+1)}\}$.
@@
Proof 4.8
Proof. Let $\mathbbm{1}(i\in S^{r}_{j})$ be an indicator random variable that is 1 if $i\in S^{r}_{j}$ and 0 otherwise. Using weak-rationality of $\phi$ (Assumption 2), we have,
$\mathbb{E}[R(S^{r}_{j})]=\sum_{i\in S^{*}_{j}}r_{i}\mathbb{E}[\mathbbm{1}(i\in S^{r}_{j})\,\phi(i,S^{r}_{j})]\geq\sum_{i\in S^{*}_{j}}r_{i}\mathbb{E}[\mathbbm{1}(i\in S^{r}_{j})]\phi(i,S^{*}_{j}).$
We claim that, $\mathbb{E}[\mathbbm{1}(i\in S^{r}_{j})]=1-\left(1-\frac{1}{j}\right)^{j}$. Note that $\mathbb{E}[1-\mathbbm{1}(i\in S^{r}_{j})]$ is the probability that $i$ is never sampled when we independently sample (with replacement) $j$ products from $S^{*}_{j}$. Observe that,
$\mathbb{E}[1-\mathbbm{1}(i\in S^{r}_{j})]=\left(1-\frac{1}{j}\right)^{j},$
here $1-\frac{1}{j}$ is the probability that a randomly sampled product from $S^{*}_{j}$ is not $i$. Note that we have equality due to the fact that $|S^{*}_{j}|=j$ (as discussed in Section 2.1). Finally, note that
$1-\left(1-\frac{1}{j}\right)^{j}\geq 1-\frac{1}{e}\qquad\forall j\geq 1$
[…]
@@
For a line graph, $\sum_{\ell=1}^{\log_{2}(\rho)}\mathbb{E}[\mathcal{R}(X^{r}_{2^{\ell-1}j_{s}})]\geq\alpha(1-1/e)\textsc{OPT}$.
@@
Proof.
The proof mimics the main steps in the proof of Lemma 4.3. The main difference is that we lose a constant factor in the guarantee since some products may be repeated due to randomized placement (Lemma 4.7). Let $O:G\to N$ denote an optimal placement, i.e., $\mathcal{R}(O)=\textsc{OPT}$. Let $O(j_{1}:j_{2})$ denote the set of products placed in locations $\{j_{1},j_{1}+2,\cdots,j_{2}\}$. For any given $\ell\in[\log_{2}(\rho)]$, consider the placement solution where we place products $O(2^{\ell-1}j_{s}:2^{\ell}j_{s}-1)$ between $2^{\ell-1}j_{s}$ and $12^{\ell}j_{s}-$ 1, same as $O$, and keep all other positions vacant. Let $O_{\ell}$ denote this solution. From the proof of Lemma 4.3, we have,
$\mathcal{R}(O)\leq\sum_{\ell=1}^{\log_{2}(\rho)}\mathcal{R}(O_{\ell}).$
Consider some $\ell\in[\log_{2}(\rho)]$. As in Lemma 4.3, we define the subset of locations $G_{\ell,k}$ as
$G_{\ell,k}=(2^{\ell-1}j_{s}:\min\{k,2^{\ell}j_{s}-1\}).$
The total revenue of random placement $X^{r}_{2^{\ell-1}j_{s}}$ can be lower bounded by the revenue of placement $O_{\ell}$ as follows,
$\mathcal{R}(O_{\ell})$ $=$ $\sum_{j=2^{\ell-1}j_{s}}^{2^{\ell}j_{s}-1}r_{O(j)}\sum_{k=j}^{j_{l}}\theta_{k}\phi\left(O(j),O(G_{\ell,k})\right),$
$=$ $\sum_{k=2^{\ell-1}j_{s}}^{j_{l}}\theta_{k}\,R(O(G_{\ell,k})),$
$\leq$ $\frac{1}{\alpha}\sum_{k=2^{\ell-1}j_{s}}^{j_{l}}\theta_{k}\,R\left(S^{*}_{2^{\ell-1}j_{s}}\right),$
$\leq$ $\frac{1}{\alpha(1-1/e)}\sum_{k=2^{\ell-1}j_{s}}^{j_{l}}\theta_{k}\,\mathbb{E}[R\left(S^{r}_{2^{\ell-1}j_{s}}\right)],$
$=$ $\frac{1}{\alpha(1-1/e)}\mathbb{E}[\mathcal{R}(X^{r}_{2^{\ell-1}j_{s}})],$
where the first inequality follows from (1) and the second inequality follows from Lemma 4.7. Thus,
$\alpha(1-1/e)\mathcal{R}(O)\leq\sum_{\ell=1}^{\log_{2}(\rho)}\mathbb{E}[\mathcal{R}(X^{r}_{2^{\ell-1}j_{s}})].$
[…]
Using the fact that,
$\sum_{\ell=1}^{\log_{2}(\rho)}\mathbb{E}[\mathcal{R}(X^{r}_{2^{\ell-1}j_{s}})]\leq(\log_{2}\rho)E[\mathcal{R}(Y)],$
we have that $Y$ is a $\alpha\frac{(1-1/e)}{\log_{2}\rho}$-approximation for a line graph as a direct corollary of Lemma 4.9.
Proof 4.11
Proof of Theorem 4.6. Fix an arbitrary permutation $\pi:[m]\to[m]$ and let $W^{\pi}(X)$ denote the expected total revenue from placement $X$ when customers visit a subset of locations $\{\pi(1),\pi(2),\cdots,\pi(m)\}$ in order, i.e.,
@@
Let $W^{\pi}(O)$ denote the expected total revenue of the optimal placement when customers visit locations in the order given by $\pi$. Note that fixing $\pi$ reduces the problem to an instance of Placement on a Line. From Lemma 4.9, we have
$\sum_{\ell=1}^{\log_{2}(\rho)}\mathbb{E}[W^{\pi}(X^{r}_{2^{\ell-1}j_{s}})]\geq\alpha(1-1/e)W^{\pi}(O).$
Now, the total revenue of placement $X$ can be written as,
@@
$(\log_{2}\rho)\,\mathbb{E}[\mathcal{R}(Y)]$ $=$ $(\log_{2}\rho)\,\sum_{\pi\in\Pi}\mathbb{E}[W^{\pi}(Y)],$
$\geq$ $\sum_{\pi\in\Pi}\sum_{\ell=1}^{\log_{2}\rho}\mathbb{E}[W^{\pi}(X^{r}_{2^{\ell-1}j_{s}})],$
$\geq$ $\sum_{\pi\in\Pi}\alpha(1-1/e)\,W^{\pi}(O),$
$=$ $\alpha(1-1/e)\textsc{OPT}.$

### v4

@@ 3.2 Hardness of Placement
Recall that the cardinality constrained assortment optimization problem is a special case of Placement (Section 2.1). In this subsection, we argue that, unless P=NP, a (11/e)(1-1/e) approximation is the best achievable guarantee in polynomial time for cardinality constrained assortment optimization under a Mixture of Multinomial Logit choice model (MMNL) and uniformly priced products. This proves that the (11/e)(1-1/e)-approximation guarantee for the Placement problem with uniformly priced products (derived in the previous section) is best possible.
@@
The proof of this theorem is included in Appendix 8.
@@ 4 Heterogeneous Prices
For heterogeneous prices, we show the following hardness result for Placement on a Line.
Lemma 4.1
Placement on a Line is weakly NP-hard [^2] A problem is weakly NP-hard if it is NP-hard, but a pseudo-polynomial time algorithm may exist for its solution. even when the choice model $\phi$ is MNL.
We show this using the hardness of product framing problem shown by Gallego et al. (2020). The proof is included in Appendix 9. Thus, unless P=NP, it is not possible to obtain a polynomial time algorithm for solving the Placement problem exactly even on a line and for the MNL choice model.
@@ 4.1 Warm-up: Placement on a Line
The proof of Lemma 4.2 is included in Appendix 9. From Lemma 4.2, we observe that selecting products solely from a fixed $S_{k}^{*}$ can be highly restrictive. In the worst case, the expected revenue from such a placement does not exceed $\frac{O(1)}{\sqrt{m}}\cdot\textsc{OPT}$. Motivated by this lemma, we explore algorithms that, for each $k\in[m]$, compute a placement using products from $S_{k}^{*}$ and then select the best among these placements. We demonstrate that this approach markedly enhances the solution to the Placement on a Line problem.
Importance of product selection from the Best of Many $\bf S_{k}^{*}$. For each $k\in[m]$, let $X^{*}_{k}$ represent the placement that positions products from $S^{*}_{k}$ in the first $k$ locations of the line (in any given order) while leaving the subsequent locations empty. Define $Y$ as the most optimal placement from the set $\{X^{*}_{k}\}_{k\in[m]}$. Specifically, consider the placement solution for the problem given by:
@@
In the next lemma, we establish that $Y$ gives $\frac{\alpha}{\log\rho}$ approximation to Placement on a Line. The result holds for all choice models satisfying Assumption 2, including multi-purchase models. Recall from Remark 2.2, that we can always fill empty locations with the highest price product without decreasing revenue.
Recall that, $j_{\min}$ ($\geq 1$) and $j_{\max}$ denote the size of the smallest (non-empty) set and largest set (respectively) with non-zero support in the browsing distribution. Also, $\rho=\frac{j_{\max}+1}{j_{\min}}=2^{d},$ for some integer $d\in\{1,2,\cdots,\lceil\log_{2}(m+1)\rceil\}$.
@@
Let $O:G\to N$ denote an optimal placement for Placement on a Line and let $\mathcal{R}(O)=\textsc{OPT}$. Recall that $O(1:k)$ denotes the set of products placed in locations $\{1,\cdots,k\}$. For $\ell\geq 1$, let $\mathcal{R}_{\ell}(O)$ denote the revenue contribution from products in location $2^{\ell-1}j_{\min}$ to $2^{\ell}j_{\min}-1$, i.e.,
$\mathcal{R}_{\ell}(O)=\sum_{j=2^{\ell-1}j_{\min}}^{2^{\ell}j_{\min}-1}\left[r_{O(j)}\,\sum_{k\geq j}\theta_{k}\,\phi\left(O(j),O(1:k)\right)\right].$ (2)
@@
We have for any $\ell\in[d]$, $(2^{\ell-1}j_{\min}:\min\{k,\,2^{\ell}j_{\min}-1\})\subseteq(j_{\min}:j_{\max}).$ To simplify notation, denote the subset of locations
$G_{\ell,k}=(2^{\ell-1}j_{\min}:\min\{k,\,2^{\ell}j_{\min}-1\}).$
We have $G_{\ell,k}\subseteq(2^{\ell-1}j_{\min}:2^{\ell}j_{\min}-1)$. Therefore, by using (2) and the weak-rationality of $\phi$ (Assumption 2), we get
$\mathcal{R}_{\ell}(O)\leq\sum_{j=2^{\ell-1}j_{\min}}^{2^{\ell}j_{\min}-1}r_{O(j)}\sum_{k=j}^{j_{\max}}\theta_{k}\phi\left(O(j),O(G_{\ell,k})\right).$
Observe that the RHS is the expected total revenue of the placement solution where we place products $O(2^{\ell-1}j_{\min}:2^{\ell}j_{\min}-1)$ between $2^{\ell-1}j_{\min}$ and $2^{\ell}j_{\min}-1$, same as $O$, and keep all other positions vacant. Let $O_{\ell}$ denote this solution. We have
Consider some $\ell$ in $[d]$. The total revenue of placement $X^{*}_{2^{\ell-1}j_{\min}}$ can be lower bounded by the revenue of placement $O_{\ell}$ as follows,
$\mathcal{R}(O_{\ell})$ $=$ $\sum_{j=2^{\ell-1}j_{\min}}^{2^{\ell}j_{\min}-1}r_{O(j)}\sum_{k=j}^{j_{\max}}\theta_{k}\phi\left(O(j),O(G_{\ell,k})\right),$
$=$ $\sum_{k=2^{\ell-1}j_{\min}}^{j_{\max}}\theta_{k}\,R(O(G_{\ell,k})),$
$\leq$ $\frac{1}{\alpha}\sum_{k=2^{\ell-1}j_{\min}}^{j_{\max}}\theta_{k}\,R\left(S^{*}_{2^{\ell-1}j_{\min}}\right),$
$\leq$ $\frac{1}{\alpha}\mathcal{R}(X^{*}_{2^{\ell-1}j_{\min}}),$
where the first inequality follows from (1) and the fact that $|O(G_{\ell,k})|\leq|G_{\ell,k}|\leq 2^{\ell-1}j_{\min}$. Thus,
$\mathcal{R}(O)\leq\frac{1}{\alpha}\sum_{\ell=1}^{d}\mathcal{R}(X^{*}_{2^{\ell-1}j_{\min}})\leq\frac{\log_{2}(\rho)}{\alpha}\,\mathcal{R}(Y),$
@@
Remark 4.5
Recall that in the proof of Lemma 4.3, we make use of at most $\log_{2}\rho$ candidate placements. In Section 4.3, we show that our analysis is tight even if we take the best placement out of the $m$ candidates $\{X^{*}_{k}\}_{k\in[m]}$. Different methods for placing/ordering the products from sets $S^{*}_{k}$ do not improve the worst case guarantee.
@@ 4.2 Main Results: Approximation Guarantee
Given an assortment, finding the optimal placement for general browsing distributions is significantly more involved as customers may start browsing from various locations and may also travel between locations in myriad ways. In particular, the optimal solution may now require copies of a product to be placed in multiple locations. We illustrate the importance of repeating products with a simple example. Recall that, for Placement on a Line the optimal placement without product repetition has the same revenue as the optimal placement with product repetition (see Remark 2.4).
@@
denote the final (randomized) placement solution. We can compute $Y$ by finding the highest revenue placement from the set $\{X^{r}_{k}\}_{k\in[m]}$ of $m$ candidate solutions. For any given $k\in[m]$, the random placement $X^{r}_{k}$ is oblivious to how customers travel between locations. In contrast, for the Placement on a Line problem, we greedily placed the products in $S^{*}_{k}$ at the top $k$ spots. The oblivious nature of randomized placement is pivotal in proving analytical guarantees.
We use a method akin to the probabilistic method to prove a guarantee on the expected performance of the algorithm. Specifically, we will show a lower bound on the expected value $\mathbb{E}[\mathcal{R}(Y)]$ of the randomized solution $Y$. Standard repetition can be used to obtain an algorithm that generates a good placement with high probability. Later, we also give a derandomized algorithm for the Markov chain choice model. Let OPT represent the optimal value of the Placement problem. Recall that, $\rho=\frac{j_{\max}+1}{j_{\min}}=2^{d}$ for some integer $d\in\{1,2,\cdots,{\log_{2}(m+1)}\}$. We establish the following performance guarantee for our algorithm.
@@
To prove Theorem 4.7, we first show the result for a line graph (Lemma 4.9). We then use the result for a line graph as a building block to establish Theorem 4.7. We start by describing the main ideas at a high level.
Proving the result for a line: Recall that in Lemma 4.3 we showed that the best (deterministic) placement from the set $\{X^{*}_{k}\}_{k\in[m]}$, is at least $\frac{\alpha}{{\log_{2}\rho}}$-approximate. Also, recall that the exact order of placement had no impact on the result, i.e. $X^{*}_{k}$ can be any placement where the products in $S^{*}_{k}$ are placed in the first $k$ spots on the line. In the new algorithm, we allow repetition of products and the candidate placement $X^{r}_{k}$ may not place some products in $S^{*}_{k}$ at any location. In fact, the value of $\mathcal{R}(X^{r}_{k})$ may be very low with some probability, such as, when a “low revenue” product is placed at every location. In Lemma 4.8, we show that such possibilities have a limited effect on the expected revenue. Then, in Lemma 4.9, we show that the sum of the expected revenues of the $\log_{2}(\rho)$ candidate placements $\{X^{r}_{j_{\min}},X^{r}_{2j_{\min}},\cdots,X^{r}_{2^{d-1}j_{\min}}\}$, is within a constant factor of the optimal solution.
@@
For a line graph, $\sum_{\ell=1}^{\log_{2}(\rho)}\mathbb{E}[\mathcal{R}(X^{r}_{2^{\ell-1}j_{\min}})]\geq\alpha(1-1/e)\textsc{OPT}$.
Using the fact that,
$\sum_{\ell=1}^{\log_{2}(\rho)}\mathbb{E}[\mathcal{R}(X^{r}_{2^{\ell-1}j_{\min}})]\leq(\log_{2}\rho)E[\mathcal{R}(Y)],$
we have that $Y$ is a $\alpha\frac{(1-1/e)}{\log_{2}\rho}$-approximation for a line graph as a direct corollary of Lemma 4.9. The proofs of Lemma 4.8 and 4.9 are included in Appendix 9.
From a line to the general case: At a high level, our approach is based on the fact that the browsing distribution of a general instance can be ’decomposed’ into a weighted ’sum’ of browsing distributions on line graphs. To make this precise, we introduce an equivalent (though perhaps less intuitive) characterization of the browsing distribution. Let $\pi:[m]\to[m]$ denote a permutation over $[m]$ and let $\Pi$ denote the set of all such permutations. For a permutation $\pi$, define $\theta^{\pi}_{j}$ as the probability that customer visits locations $\{\pi(1),\pi(2),\cdots,\pi(j)\}$ exactly in this order [^3] Let $L=\{\pi(1),\pi(2),\cdots,\pi(j)\}$. Recall that, $\mathbb{P}_{B}(L)$ represents the probability that $L$ is the set of locations visited by the customer in any order. In contrast, $\theta^{\pi}_{j}$ is the probability that the customer visits the locations in the specified order.. Let $X^{\pi}(1:j)$ denote the set of products placed at $\{\pi(1),\cdots,\pi(j)\}$. Then the total expected revenue can be expressed as a sum over all permutations,
Now fix a permutation $\pi\in\Pi$. Consider the line graph with ordered set of locations $\{\pi(1),\cdots,\pi(m)\}$, where the customer visits the first $j$ locations with probability $\theta^{\pi}_{j}$, and visits none with probability $1-\sum_{j\in[m]}\theta^{\pi}_{j}$. Using Lemma 4.9 on this line graph, we have that the sum of the expected revenues of $\log_{2}(\rho)$ (randomized) candidate placements is a constant factor of the optimal revenue. Since the candidate solutions constructed are independent of the specific permutation $\pi$, combining this result with the decomposition above yields Theorem 4.7.
@@
Proof of Theorem 4.7. Fix an arbitrary permutation $\pi\in\Pi$ and let $W^{\pi}(X)$ denote the expected total revenue from placement $X$ when customers visit a subset of locations $\{\pi(1),\pi(2),\cdots,\pi(m)\}$ in order, i.e.,
@@
Let $W^{\pi}(O)$ denote the expected total revenue of the optimal placement when customers visit locations in the order given by $\pi$. Note that fixing $\pi$ reduces the problem to an instance of Placement on a Line where, for every $j\in[m]$, the customer visits the first $j$ locations with probability $\theta^{\pi}_{j}$, and visits none of the locations with probability $1-\sum_{j\in[m]}\theta^{\pi}_{j}$. From Lemma 4.9, we have
$\sum_{\ell=1}^{\log_{2}(\rho)}\mathbb{E}[W^{\pi}(X^{r}_{2^{\ell-1}j_{\min}})]\geq\alpha(1-1/e)W^{\pi}(O),$
here $W^{\pi}(O)$ is a lower bound on the optimal revenue for this instance of Placement on a Line. Now, the total revenue of placement $X$ can be written as,
@@
$(\log_{2}\rho)\,\mathbb{E}[\mathcal{R}(Y)]=(\log_{2}\rho)\,\sum_{\pi\in\Pi}\mathbb{E}[W^{\pi}(Y)]$ $\geq$ $\sum_{\pi\in\Pi}\sum_{\ell=1}^{\log_{2}\rho}\mathbb{E}[W^{\pi}(X^{r}_{2^{\ell-1}j_{\min}})],$
$\geq$ $\sum_{\pi\in\Pi}\alpha(1-1/e)\,W^{\pi}(O)=\alpha(1-1/e)\textsc{OPT}.$